### v3

@@ 3 Graphs with Polynomial Number of Steiner Trees
Figure 2: Overview of graph classes and instances for which the Requirement Cut problem admits an improved (single log) approximation in our paper.
@@ 4.1 Linear Program
Summary of the proof: Starting from any $S$-Steiner tree whose cuts in $G$ do not split $S$ into $r$ terminal components, $T_{0}$. We successively find other trees that must have been considered by the LP and were also cut enough times. These other trees form a family of $S$-Steiner trees that use 0-cost edges and the union of their cut-edges in $G$ partitions $S$ into at least $r$ components.
@@ 5 Depth of Series Parallel Graphs
In this section, we refine our approximation bounds for the Requirement Cut problem on series-parallel graphs by analyzing differently the algorithm given by Emek and Peleg [emek_tight_2010]. This different analysis coupled with the $\mathcal{O}(\log(g))$-approximation for trees [gupta_improved_2010] yields a $\mathcal{O}(m\cdot\log(g))$-approximation for series-parallel graphs of depth $m$. Series-parallel graphs in general cannot be embedded into trees with less than $\log(n)$ expected distortion [gupta_cuts_2004]. This means that general series-parallel graphs currently have the same approximation ratio for our problem as general graphs, i.e., $\mathcal{O}(\log(g)\cdot\log(n))$. In fact, the infinite class of series-parallel graphs that Gupta et al. [gupta_cuts_2004] introduce as having distortion no less than $\log(n)$ when embedded into trees have a 0ptof $\mathcal{O}(\log(n))$. Meaning that our result provides a credible alternative also in some cases when the 0ptis smaller than $\log(n)$ and not only when it is a given constant.

### v4

@@ 3 Graphs with Polynomial Number of Steiner Trees
Figure 2: Overview of graph classes and instances for which the Requirement Cut problem admits a single-logarithmic approximation in this paper.
@@ 5 Depth of Series Parallel Graphs
In this section we refine the approximation bounds for the Requirement Cut problem on series-parallel graphs by reanalyzing the embedding algorithm of Emek and Peleg [emek_tight_2010]. Combined with the $\mathcal{O}(\log g)$-approximation for trees [gupta_improved_2010], this yields an $\mathcal{O}(m\log g)$-approximation on series-parallel graphs of depth $m$ (as in Definition 2.8).Series-parallel graphs in general cannot be embedded into trees with expected distortion $o(\log n)$ [gupta_cuts_2004]. Hence, for unrestricted series-parallel graphs, the best known ratio for Requirement Cut matches the general case, namely $\mathcal{O}(\log g\cdot\log n)$ via tree embeddings [nagarajan_approximation_2010, gupta_improved_2010]. The lower-bound construction of Gupta et al. [gupta_cuts_2004] in fact has depth $\Theta(\log n)$, so our result gives a strictly better guarantee whenever the depth $m$ is $o(\log n)$ (and in particular when $m$ is constant).